A home-made lefsa enterprise estimates that they can sell 30 dozen lefsa at $3.00 per dozen and 25 dozen lefsa at $4.00 per dozen. write a linear equation for the amount of lefsa sold in terms of the price per dozen. use your equation to approximate the amount of lefsa sold at $5.00 per dozen.
thanks for helping me on this!!

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

A home-made lefsa enterprise estimates that they can sell 30 dozen lefsa at $3.00 per dozen and 25 dozen lefsa at $4.00 per dozen. write a linear equation for the amount of lefsa sold in terms of the price per dozen. use your equation to approximate the amount of lefsa sold at $5.00 per dozen.
-----------------------------------
You have two points of the form (price,lefsa)
They have (3,30) and (4,25)
-------------------------------
m = (25-30)/(4-3) = -5
---------------------------
You have m=-5 and y = 25 when x = 4; solve for b in the form y = mx + b
-------------
25 = -5*4 + b
b = 45
----------
lefsa(in dozens) = -5(price) + 45
lefsa($5) = -5*5 + 45
lefsa($5) = 20 dozen
========================
Cheers,
Stan H.

Answer by solver91311(24713)



Here you have a situation where the value 3 maps to the value 30 and the value 4 maps to the value 25. This is exactly the same as saying that you have two ordered pairs defining a straight line and the two ordered pairs are (3, 30) and (4, 25).

Notice that when the x value increases by 1, the y value decreases by 5. That means that the slope of our line is -5. Also notice, using this slope number, that 2 would map to 35, 1 would map to 40, and 0 would map to 45. That tells us that the y-intercept of the line is 45.

Now we can write the equation in slope intercept form directly:



To calculate the sales at a $5 price point, substitute 5 for x and do the arithmetic.


John